please help me with this problem
complete the following table for the equation y=6x-5
[x][-3][-1][ ][ ]
[y][ ][ ][2][4]

Answer by uma(370)

The given equation is y = 6x - 5
When x = -3,
y = 6(-3)-5
= - 18 - 5
= - 23
When x = -1,
y = 6(-1)-5
= -6 - 5
= - 11
When y = 2,
6x - 5 = 2
==> 6x -5 + 5 = 2 + 5
==> 6x = 7
==> 6x/6 = 7/6
==> x = 7/6
When y = 4,
6x - 5 = 4
==> 6x - 5 + 5 = 4 + 5
==> 6x = 9
==> 6x/6 = 9/6
==> x = 3/2
The table is as follows:
x | -3| -1| 7/6 |3/2|
__________________________
y |-23|-11|2 | 4|

Good Luck!!!
________________________